A BOUND FOR CANONICAL DIMENSION OF THE SPINOR GROUP

NIKITA A. KARPENKO

Abstract. Using the theory of non-negative intersections, duality of the Schubert varieties, and Pieri-type formula for a maximal orthogonal grassmannian, we get an upper bound for the canonical dimension cd(Spin_{n}) of the spinor group Spin_{n}. A lower bound is given by the canonical 2-dimension cd_{2}(Spin_{n}), computed in [8]. If n or n + 1 is a power of 2, no space is left between these two bounds; therefore the precise value of cd(Spin_{n}) is obtained for such n.

In the appendix, we also produce an upper bound for canonical dimension of the semi-spinor group (giving the precise value of the canonical dimension in the case when the rank of the group is a power of 2), compute canonical dimension of the projective orthogonal group, and show that the spinor group represents the unique difficulty when trying to compute the canonical dimension of an arbitrary simple split group, possessing a unique torsion prime.

Contents

1. Introduction 2 2. Non-negativity 3 3. Dual Schubert varieties 3 4. Pieri formula 5 5. Proof of Theorem 1.1 6 6. Appendix: Split simple algebraic groups of arbitrary type 6 6.1. A_{n}, n ≥ 1 7 6.2. C_{n}, n ≥ 2 7 6.3. B_{n}, n ≥ 3 7 6.4. D_{n}, n ≥ 4 7 6.5. G_{2} 9 6.6. F_{4} and E_{n}, n = 6, 7, 8 9 References 9

Date: December, 2004. Key words and phrases. Algebraic groups, projective homogeneous varieties, Chow groups.

2000 Mathematical Subject Classifications: 14L17; 14C25.

The James D. Wolfensohn Fund and The Ellentuck Fund support is acknowledged. Partially supported by the European Community’s Human Potential Programme under contract HPRN-CT-2002-00287, KTAGS.

1. Introduction

Let F be an arbitrary field (of an arbitrary characteristic). Let X be a smooth algebraic variety over F. A field L ⊃ F is called a splitting field (of X), if X(L) ≠ ∅. A splitting field L is called generic, if for any splitting field L′ there exists a place L → L′. The canonical dimension cd(X) is defined as the minimum of transcendence degrees of generic splitting fields of X (cf. [8, §2]).

Let G be an algebraic group over F. The canonical dimension cd(G) of G, as introduced in [1], is the maximum of canonical dimensions of G-torsors, defined over field extensions of F (of course, it is not the same as the canonical dimension of the underlying variety of G, which is not an interesting invariant because is always 0). For the spinor group, it is explained in loc.cit. that cd(Spin_{2n+1}) = cd(Spin_{2n+2}) (where n is an integer ≥ 1), so that we will discuss only cd(Spin_{2n+1}) here.

Although the canonical dimension of, say, a smooth projective variety X can be expressed in terms of algebraic cycles on X (see [8, cor. 4.7]), there are no general recipes for computing cd(X) or cd(G). A better situation occurs with the canonical p-dimension cd_{p}, a p-relative version of cd, where p is a prime, defined in [8, §3]: a recipe for computing cd_{p}(G) of an arbitrary split simple G is obtained in loc.cit. In particular, one has

cd_{2}(Spin_{2n+1}) = n(n + 1)/2 − 2^{r} + 1 ,

where r is the smallest integer such that 2^{r} > n (while cd_{p}(Spin_{2n+1}) = 0 for any odd prime p). Since cd(G) ≥ cd_{p}(G) for any G and p, we have a lower bound for the canonical dimension of the spinor group, given by its canonical 2-dimension:

cd(Spin_{2n+1}) ≥ \frac{n(n + 1)}{2} − 2^{r} + 1 .

In this note we establish the following upper bound for cd(Spin_{2n+1}):

Theorem 1.1. For any n ≥ 1, one has cd(Spin_{2n+1}) ≤ n(n − 1)/2.

The proof is given in section 5. It makes use of the theory of non-negative intersections, of duality between Schubert varieties, and of the Pieri formula for the split maximal orthogonal grassmannian.

Note that the lower bound for cd(Spin_{2n+1}) coincide with the upper one if (and only if) n + 1 is a power of 2. Therefore, for such n, we get the precise value:

Corollary 1.2. If n + 1 is a power of 2, then cd(Spin_{2n+1}) = n(n − 1)/2. □

Remark 1.3. For n up to 4, it is easy to see that cd(Spin_{2n+1}) = cd_{2}(Spin_{2n+1}) (see [1]), but for every n ≥ 5 the bound of Theorem 1.1 is new.

In the last section we explain why the question about canonical dimension of the spinor group is of particular importance, produce an upper bound for canonical dimension of the semi-spinor group (see Theorem 6.2 and Corollary 6.3), and compute canonical dimension of the projective orthogonal group (see Theorem 6.1).

Acknowledgments: This work was accomplished during my stay at the Princeton Institute for Advanced Study, a perfect place to work.

2. Non-negativity

By scheme we mean a separated scheme of finite type over a field. Variety is an integral scheme.

Let X be a scheme. Following [4, ch. 12], an algebraic cycle n_{1}Y_{1}+… n_{r}Y_{r} on X (where r is ≥ 0, Y_{i} are closed subvarieties of X, and n_{i} are integers) is called non-negative, if the coefficients n_{1}, …, n_{r} are non-negative. An element α of the integral Chow group CH(X) is called non-negative, if it can be represented by a non-negative cycle.

I thank I. Panin for pointing me out the following fact:

Lemma 2.1. Let L be a line vector bundle over a smooth variety X. The first Chern class c_{1}(L) ∈ CH^{1}(X) is non-negative if and only if L has a non-zero global section.

Proof. Fixing an imbedding of L into the constant O_{X}-module F (X)^{∗}, we get a 1codimensional cycle C on X, representing c_{1}(L), such that a function f ∈ F (X)^{∗} is a global section of L if and only if the cycle div(f)+C is non-negative, [5]. Since C is rationally equivalent to a non-negative cycle if and only if the cycle div(f) + C is non-negative for some f ∈ F (X)^{∗}, we are done. □

Corollary 2.2. Let X be a smooth absolutely irreducible variety over a field F, α an element of CH^{1}(X), E/F a field extension. If α_{E} ∈ CH(X_{E}) is non-negative, then α itself is non-negative.

Proof. Let L be a line vector bundle over X such that c_{1}(X) = α. Assume that α_{E} is non-negative. Then the E-vector space Γ(X_{E}, L_{E}) of global sections of L_{E} is non-zero by Lemma 2.1. Since Γ(X_{E}, L_{E}) = Γ(X, L) ⊗_{F} E, it follows that Γ(X, L) ≠ 0; therefore, once again by Lemma 2.1, α is non-negative. □

We are going to use the following

Theorem 2.3 ([4, §12.2]). Let X be a smooth variety such that its tangent bundle is generated by the global sections. Then the product of non-negative elements in CH(X) is non-negative. Moreover, if α ∈ CH(X) is represented by a non-negative cycle with support A ⊂ X, while β ∈ CH(X) is represented by a non-negative cycle with support B ⊂ X, then the product αβ ∈ CH(X) can be represented by a non-negative cycle with support on the intersection A ∩ B.

Remark 2.4. If X is a projective homogeneous variety under an action of an algebraic group, then the tangent bundle of X is generated by the global sections. Indeed, there exists a field extension E/F such that the variety X_{E} is isomorphic to the quotient G/P of a semisimple algebraic group G over E modulo a parabolic subgroup P ⊂ G. Therefore the tangent bundle of the variety X_{E} is generated by the global sections. Since the property of being generated by global sections is not changed under extension of the base field, the tangent bundle of the variety X is also generated by the global sections.

3. Dual Schubert varieties

Let G be a split semisimple algebraic group, T ⊂ B ⊂ G a maximal split torus and a Borel subgroup of G. Let W be the Weyl group of G, and let S ⊂ W be the set of reflections with respect to the simple roots. We fix a subset θ ⊂ S, take the subgroup W_{θ}, generated by θ, and consider the parabolic subgroup P = P_{θ} = BW_{θ}B ⊂ G.

Using the length function l: W → ℤ_{≥0}, induced by the set S of generators of the group W, we take in each coset of W/W_{θ} the unique smallest length element and write W^{θ} ⊂ W for the set of representatives thus obtained.

The variety X = G/P is cellular, the cells BwP/P are indexed by w ∈ W^{θ}. We write X_{w} for the closure in G/P of the corresponding cell. The varieties X_{w} are called (generalized) Schubert varieties; their classes [X_{w}] ∈ CH(X), called (generalized) Schubert classes, form a basis of the group CH(X). Moreover, dim X_{w} = l(w).

Sometimes the upper indexation of the Schubert varieties, which respects their codimension, is more convenient. To define it, let us take the (unique) largest length element w_{0} ∈ W and set X^{w} = X_{w′}, where w′ is the smallest length element in the coset w_{0}wW_{θ} (so that w′ = w_{0}w′′, where w′′ is the largest length element of the coset wW_{θ}). Now we have codim X^{w} = dim X_{w} = l(w).

Proposition 3.1 ([10, prop. 1.4]). Let deg : CH(X) → ℤ be the degree homomorphism. Then for any w, w′ ∈ W^{θ} one has:

deg([X^{w}] · [X_{w′}]) = \begin{cases}1, if w′ = w; \\ 0, otherwise. \end{cases}

Because of this property, we refer to the varieties X^{w} and X_{w} (as well as to their classes) as mutually dual.

Let us now specify the situation: take as G the special orthogonal group SO(ϕ) of a split quadratic form ϕ: V → F, where V is a (2n+1)-dimensional vector space over a field F. Saying split, we mean existence in V of an n-dimensional totally isotropic subspace. Let us choose a complete flag

F = (0 = F_{0} ⊂ F_{1} ⊂ ⋯ ⊂ F_{n}) , dim F_{i} = i

of totally isotropic subspaces of (V, ϕ) and a subspace D_{i} in each F_{i} such that F_{i} = F_{i−1} ⊕ D_{i}. Then we take as B ⊂ G the stabilizer of F, as P the stabilizer of F_{n}, and as T the stabilizer of all D_{i}. The variety X = G/P is therefore the maximal orthogonal grassmannian of ϕ.

Now the Schubert varieties on X are indexed by the strictly decreasing sequences i_{1} > i_{2} > ⋯ > i_{s} of positive integers, satisfying n ≥ i_{1}. The (i_{1} … i_{s})-th Schubert variety X^{i_{1}...i_{s}} is the closed subvariety of the subspaces W ∈ X such that dim(W ∩ F_{n+1−i_{t}}) ≥ t for t = 1, 2, …, s; the variety X^{i_{1}...i_{s}} has codimension i_{1} + ⋯ + i_{s}, and we write e_{i_{1}...i_{s}} for its class in CH(X). The Schubert classes e_{i} (i = 1, 2, …, n) are called special.

As a specific case of Proposition 3.1 we get

Corollary 3.2. For any e_{i_{1}...i_{s}} there exists e_{i′_{1}...i′_{s′}} such that for any e_{i′′_{1}...i′′_{s′′}}

deg(e_{i′′_{1}...i′′_{s′′}} · e_{i′_{1}...i′_{s′}}) = \begin{cases}1, if e_{i′′_{1}...i′′_{s′′}} = e_{i_{1}...i_{s}}; \\ 0, otherwise. \end{cases}

□

4. Pieri formula

The classical Pieri formula expresses the product of a special Schubert class by an arbitrary Schubert class as a linear combination of Schubert classes in the case of a usual grassmannian.

We are going to use an analogues formula for a split maximal orthogonal grassmanian:

Theorem 4.1 ([6], see also [13]). Let n be a positive integer, X the maximal orthogonal grassmannian of a split (2n+1)-dimensional quadratic form. The following multiplication formula holds for the Schubert classes in CH(X): for any strictly decreasing sequence of positive integers x = (x_{1}, …, x_{k}), satisfying x_{1} ≤ n, and any positive integer p ≤ n, one has

e_{p} · e_{x} = ∑_{y} 2^{m_{y}^{x,p}}e_{y} ,

where the sum runs over all strictly decreasing sequences of integers y = (y_{1}, …, y_{k+1}), satisfying

n ≥ y_{1} ≥ x_{1} ≥ y_{2} ≥ ⋯ ≥ y_{k} ≥ x_{k} ≥ y_{k+1} ≥ 0

and y_{1} + ⋯ + y_{k+1} = p + x_{1} + ⋯ + x_{k} (in the case of y_{k+1} = 0, we define e_{y} as e_{y_{1}...y_{k}}). The exponent m_{y}^{x,p} of the coefficient of e_{y} is determined as follows:

m_{y}^{x,p} = \begin{cases}the quantity of i ∈ [1, k] such that y_{i} > x_{i} > y_{i+1}, & if y_{k+1} ≠ 0; \\ the above quantity minus 1, & if y_{k+1} = 0. \end{cases}

Remark 4.2. Theorem 4.1 is proved in [6] under the assumption that the base field is algebraically closed; it is proved in [13] under the assumption that the base field is ℂ. However, as shown in [3], the multiplication table for the Schubert classes in CH(G/B), where G is a split semisimple algebraic group and B is its Borel subgroup, depends only on the type of G and does not depend on the base field. Now if P ⊂ G is a parabolic subgroup, containing B, the pull-back with respect to the projection G/B → G/P is an injective ring homomorphism, mapping each Schubert class [X^{w}] ∈ CH(G/P) (the notation is introduced in §3) to the “same” Schubert class [X^{w}] ∈ CH(G/B) (see [10, lemma 1.2(b)] and [3, §3.3]).^{1} Therefore the multiplication table for the Schubert classes in CH(G/P) depends only on the type of the pair (G, P) and does not depend on the base field either.

Corollary 4.3. Under condition of Theorem 4.1, one has e_{1}^{n} = e_{n} +…, where dots stand for a linear combination of Schubert classes different from e_{n}.

Proof. Let x and y be two strictly decreasing sequences of positive integers ≤ n. Let us say that y is a deformation of x (and write x ⇝ y), if e_{y} appears in the formula for e_{1} · e_{x}, given by Theorem 4.1, in which case we refer to the number m_{y}^{x,1} as the exponent of the deformation.

There is a unique chain of deformations, transforming (1) to (n), namely, the chain (1) ⇝ (2) ⇝ … ⇝ (n). Since the exponent of each deformation in the chain is 0, the statement follows. □

1and, as follows from [10, lemma 1.2(c)], under the push-forward homomorphism, [X_{w}] is mapped to [X_{w}] for w ∈ W^{θ}, and [X_{w}] is mapped to 0 for w ∉ W^{θ}.

Remark 4.4. Unfortunately none of the Schubert classes appears with coefficient 1 in the decomposition of e^{n+1}.

5. Proof of Theorem 1.1

Let V be a (2n + 1)-dimensional vector space over F equipped with a non-degenerate quadratic form ϕ : V → F. Let X be the orthogonal grassmannian of n-dimensional (maximal) totally isotropic subspaces in V.

By [8, cor. 4.7], the canonical dimension cd(X) is the minimum of dim Y, where Y runs over all closed subvarieties of X such that Y (F (X)) ≠ ∅. As explained in [1], cd(Spin_{2n+1}) is the maximum of cd(X), taken over all ϕ such that the even Clifford algebra C_{0}(ϕ) is split. Therefore, in order to prove Theorem 1.1, it suffices to prove the following

Proposition 5.1. Let X be the maximal orthogonal grassmannian of a (2n+1)-dimensional quadratic form ϕ such that [C_{0}(ϕ)] = 0. Then there exists a closed subvariety Y ⊂ X of dimension n(n − 1)/2 such that Y (F (X)) ≠ ∅.

Proof. Since over the function field F (X) the quadratic form ϕ is split, we may speak about Schubert classes in CH(X̄), where X̄ stands for X_{F (X)}. We write C̄H(X) for the image of the restriction homomorphism CH(X) → CH(X̄).

Let us take the Schubert class e_{1} ∈ CH^{1}(X̄). By our assumption on C_{0}(ϕ), we have C̄H^{1}(X) = CH^{1}(X̄) (see, e.g., [8, §8.2] or [14, proof of lemma 3.1]; cf. proof of Theorem 6.2); therefore e_{1} ∈ C̄H^{1}(X). Moreover, by Corollary 2.2, e_{1} is non-negative. It follows by Theorem 2.3 with Remark 2.4 that the n-th power of e_{1} is also non-negative, so that we can write

e_{1}^{n} = n_{1}[Y_{1}] + ⋯ + n_{r}[Y_{r}]

with some non-negative integers n_{i} and some closed subvarieties Y_{i} ⊂ X. Note that dim Y_{i} = dim X − n = n(n − 1)/2 for all i.

By Corollary 4.3, we have e_{1}^{n} = e_{n} + ⋯ ∈ CH(X̄), where dots stand for a linear combination of Schubert classes different from e_{n}. Using Corollary 3.2, we find the Schubert class e ∈ CH(X̄) dual to e_{n}, that is, such that deg(e_{n}e) = 1 while deg(e′e) = 0 for any Schubert class e′ different from e_{n}. For this e we have deg(e_{1}^{n}e) = 1. Since the product e_{1}^{n}e is non-negative (by Theorem 2.3 with Remark 2.4), it follows that deg([Ȳ_{i}] · e) = 1 (and n_{i} = 1) for some i ∈ [1, r], where Ȳ_{i} = (Y_{i})_{F (X)}

Let Z be the Schubert variety, representing e. Since the product [Ȳ_{i}] · [Z] is a 0-cycle class of degree 1 and can be represented by a non-negative cycle on the intersection Ȳ_{i} ∩ Z (see Theorem 2.3 with Remark 2.4), the scheme Ȳ_{i} has a rational point, that is, Y_{i}(F (X)) ≠ ∅. □

6. Appendix: Split simple algebraic groups of arbitrary type

In this section we show that the spinor group represents the only difficulty when trying to answer the following question: let G be a split simple algebraic group, having a unique torsion prime p; is it true that cd(G) = cd_{p}(G)?

To explain the assumption about the uniqueness of torsion prime, made in this question, let us notice that it is not clear what to expect from cd(G) in the case when G has more than one torsion prime (recall that cd_{p}(G) > 0 if and only if p is a torsion prime of G, [8, rem. 6.10]).

Requiring the uniqueness of torsion prime p, we are in the situation where cd_{p}(G) is known (see below), so that one may expect to be able to get some interesting information also on cd(G).

Let us consider all split simple algebraic groups, having a unique torsion prime, type by type. All cited results on canonical dimension are from [1] and on canonical p-dimension from [8]. In each subsection we assume without repeating it, that G is a split simple group of the type under consideration.

6.1. A_{n}, n ≥ 1. We have G ≃ SL_{n+1} /μ_{l}, where l is a positive integer, dividing n + 1. The group G has a unique torsion prime p if and only if l is a positive power of p. In this case it is known that cd(G) = cd_{p}(G) = p^{k} − 1, where k is the largest integer such that p^{k} divides n + 1.

6.2. C_{n}, n ≥ 2. If G is simply connected, then cd(G) = 0.

If G is not simply connected, then 2 is its unique torsion prime and cd(G) = cd_{2}(G) = 2^{k+1} − 1, where k is the largest integer such that 2^{k} divides n.

6.3. B_{n}, n ≥ 3. The prime 2 is the unique torsion prime here.

If G is simply connected, then G ≃ Spin_{2n+1}.

If G is not simply connected, then G ≃ SO_{2n+1}. In this case it is known that cd(G) = cd_{2}(G) = n(n + 1)/2 (originally proved in [7]).

6.4. D_{n}, n ≥ 4. The prime 2 is the unique torsion prime here.

If G is simply connected, then G ≃ Spin_{2n}.

If G is not simply connected, then either G ≃ SO_{2n} or G ≃ PGO_{2n}^{+}. If n is odd, there are no other possibilities. If n is even, then one more possibility is added: G ≃ Spin_{2n}^{∼} (the semi-spinor group).

For G = SO_{2n}, we know that cd(G) = cd_{2}(G) = n(n − 1)/2.

In the case when G = PGO_{2n}^{+}, we know that cd_{2}(G) = n(n − 1)/2 + 2^{k} − 1, where k is the largest integer such that 2^{k} divides n. We compute cd(G) now:

Theorem 6.1. For G = PGO_{2n}^{+}, one has cd(G) = cd_{2}(G).

Proof. We only need to show that cd(G) ≤ cd_{2}(G).

The set of isomorphism classes of G-torsors is mapped surjectively and with trivial kernel onto the set of isomorphism classes of central simple algebras of degree 2n with an orthogonal pair of trivial discriminant, [9, §29F].

Let A be such an algebra. The scheme of rank n isotropic ideals in A has two (possibly non-isomorphic) components; let X be a component of this scheme. Note that dim X = n(n − 1)/2. Furthermore, let Y be the Severi-Brauer variety of A. The algebra with orthogonal pair A is split (i.e., its class is the distinguished point of the set of isomorphism classes) if and only if the variety Y ×X has a rational point. Therefore, to get the required bound for cd(G), it suffices to get the same bound for cd(Y × X) (see [8, rem. 3.7]).

The index of the algebra A_{F (X)} divides n and the index of A itself is a power of 2. Therefore, ind A_{F (X)} divides 2^{k}, and we may write A as the tensor product of a degree 2^{k} central simple algebra D and a matrix algebra M. Let Z^{o} be the Severi-Brauer variety of D; the Severi-Brauer variety of M is a projective space ℙ. Tensor product of ideals gives rise to a closed imbedding Z^{o} × ℙ ↪ Y_{F (X)} (which becomes a Segre imbedding over F (Y ×X)). Choosing a closed rational point on ℙ, we identify Z^{o} with a closed subvariety of Y_{F (X)}. Let Z be the closure of Z^{o} in Y × X.

The constructed closed subvariety Z ⊂ Y × X has the dimension

dim X + dim Z^{o} = \frac{n(n − 1)}{2} + 2^{k} − 1 ,

which coincides with the bound we want for cd(Y × X). To show that this number is really a bound for cd(Y × X), it suffices to show that Z(F (Y × X)) ≠ ∅, [8, cor. 4.7].

Let us write Z̄, X̄, and Ȳ for Z, X, and Y over the field F (Y × X). The variety Ȳ is isomorphic to a projective space; let H ⊂ Ȳ be a hyperplane. Let x ∈ X̄ be a rational point. The degree of the product [Z̄] · ([H]^{2^{k}−1} × [x]) is 1. Moreover, this product can be represented by a non-negative cycle with support on Z̄ (Theorem 2.3 with Remark 2.4). Therefore, the scheme Z̄ has indeed a rational point. □

For the semi-spinor group G = Spin_{2n}^{∼}, we know that cd_{2}(G) = n(n − 1)/2 + 2^{k} − 2^{r}, where k is the same as above, while r is the smallest integer such that 2^{r} ≥ n. We show now, how the bound for cd(Spin_{2n}) = cd(Spin_{2n−1}), established in this paper (Theorem 1.1), produces a bound for cd(Spin_{2n}^{∼}).

Theorem 6.2. For k being the largest integer such that 2^{k} divides n, one has

cd(Spin_{2n}^{∼}) ≤ \frac{(n − 1)(n − 2)}{2} + 2^{k} − 1 .

Proof. The set of isomorphism classes of G-torsors for G = Spin_{2n+2}^{∼} is mapped surjectively and with trivial kernel onto the set of isomorphism classes of degree 2n central simple algebras with an orthogonal pair of trivial discriminant and trivial component of the Clifford algebra.

Let A be such an algebra. For a component X of the scheme of rank n isotropic ideals of A, let us consider the homomorphism CH^{1}(X) → CH^{1}(XF̄), where F̄ is an algebraic closure of F. Since the absolute Galois group of F acts trivially on CH^{1}(XF̄) = ℤ · e_{1}, the cokernel of this homomorphism is identified with the relative Brauer group

Br(F (X)/F) = Ker ( Br(F) → Br(F (X)))

of the function field of X (see, e.g., [2, proof of th. 3.1]), which is trivial, if the corresponding to X component of the Clifford algebra of A is trivial, [11, appendix] (see [12, rem 9.2] for the characteristic 2 case). Therefore the homomorphism CH^{1}(X) → CH^{1}(XF̄) is surjective for an appropriate choice of X. Note that if the second component of the Clifford algebra is not trivial (what happens if and only if the algebra A is not trivial [9, (9.14)]), the component of the scheme of the ideals is uniquely determined by this condition; in particular, X can not be chosen arbitrary in this proof.

Let Y be the Severi-Brauer variety of A. The algebra A is split if and only if the variety Y × X has a rational point. Therefore, to get the required bound for cd(G), it suffices to get the same bound for cd(Y × X).

Let us agree to write T̄ for the scheme T_{F (Y ×X)}, if T is a scheme over F. The variety X̄ is isomorphic to a component of the maximal orthogonal grassmannian of a hyperbolic 2n-dimensional quadratic form and therefore is isomorphic to the maximal orthogonal grassmannian of a split (2n − 1)-dimensional quadratic form, considered in the main part of the paper. In particular, dim X = n(n − 1)/2. Moreover, C̄H^{1}(X) = CH^{1}(X̄) (where C̄H(X) stands for the image of CH(X) → CH(X̄)), so that the first special Schubert class e_{1} ∈ CH(X̄) lies in C̄H(X). Acting as in the proof of Theorem 1.1, we find a component T of the support of a non-negative cycle, representing e_{1}^{n−1}, such that deg([T̄] · e) = 1, where e ∈ CH(X̄) is the Schubert class dual to e_{n−1}. Note that dim T = (n − 1)(n − 2)/2.

Since the variety X_{F (T)} has a rational point, the index of the algebra A_{F (T)} divides n. Since the index of A itself is a power of 2, ind A_{F (T)} divides 2^{k}, and we may write A as the tensor product of a degree 2^{k} central simple algebra D and a matrix algebra M. Let Z^{o} be the Severi-Brauer variety of D. Acting as in the proof of Theorem 6.1, we identify Z^{o} with a closed subvariety of Y_{F (T)}. Let Z be the closure of Z^{o} in Y × T ⊂ Y × X.

The constructed closed subvariety Z ⊂ Y × X has the dimension

dim T + dim Z^{o} = \frac{(n − 1)(n − 2)}{2} + 2^{k} − 1 ,

which coincides with the bound we want for cd(Y × X). To show that this number is really a bound for cd(Y × X), it suffices to show that Z(F (Y × X)) ≠ ∅, i.e., that Z̄ has a rational point.

The variety Ȳ is isomorphic to a projective space; let H ⊂ Ȳ be a hyperplane. The degree of the product [Z̄]·([H^{2^{k}−1}]×e) is 1. Moreover, this product can be represented by a non-negative cycle with support on Z̄. Therefore, the scheme Z̄ has indeed a rational point. □

Note that the upper bound of Theorem 6.2 coincides with the lower bound, given by cd_{2}, if (and only if) n is a power of 2. Therefore, in this case, we get the precise value:

Corollary 6.3. If n is a power of 2, then cd(Spin_{2n}^{∼}) = n(n − 1)/2. □

6.5. G_{2}. The prime 2 is the unique torsion prime here, and one knows that

cd(G) = cd_{2}(G) = 3 .

6.6. F_{4} and E_{n}, n = 6, 7, 8. We have multiple torsion primes here.

References

[1] G. Berhuy and Z. Reichstein. On the notion of canonical dimension for algebraic groups. Adv. Math., to appear.

[2] J.-L. Colliot-Thélène, D. Madore. Surfaces de Del Pezzo sans point rationnel sur un corps de dimension cohomologique un. J. Inst. Math. Jussieu 3 (2004), no. 1, 1–16.

[3] M. Demazure. Désingularisation des variétés de Schubert généralisées. Ann. Sci. Ecole Norm. Sup. (4) 7 (1974), 53–88.

[4] W. Fulton. Intersection Theory. Springer-Verlag; Berlin, Heidelberg, New York, Tokyo, 1984.

[5] R. Hartshorne. Algebraic Geometry. Springer-Verlag; New York, Heodelberg, Berlin, 1977.

[6] H. Hiller, B. Boe. Pieri formula for SO_{2n+1}/U_{n} and Sp_{n}/U_{n}. Adv. Math. 62 (1986), 49–67.

[7] N. A. Karpenko. Canonical dimension of orthogonal groups. Transform. Groups, to appear.

[8] N. A. Karpenko and A. S. Merkurjev. Canonical p-dimension of algebraic groups. Linear Algebraic Groups and Related Structures (Preprint Server), December 2004, 22 p.

[9] M.-A. Knus, A. Merkurjev, M. Rost, J.-P. Tignol. The Book of Involultions. American Mathematical Society Colloquium publications, v. 44, 1998.

[10] B. Köck. Chow motif and higher Chow theory of G/P. Manuscripta Math. 70 (1991), 363–372.

[11] A. S. Merkurjev, I. A. Panin, A. R. Wadsworth. Index reduction formulas for twisted flag varieties, I. K-Theory 10 (1996), no. 6, 517–596.

[12] A. S. Merkurjev, I. A. Panin, A. R. Wadsworth. Index reduction formulas for twisted flag varieties, II. K-Theory 14 (1998), no. 2, 101–196.

[13] P. Pragacz, J. Ratajski. Pieri formula for isotropic Grassmannians; the operator approach. Manuscripta Math. 79 (1993), no. 2, 127–151.

[14] B. Totaro. The torsion index of the spin group. Duke Math. J., to appear.

Laboratoire de Mathématiques de Lens, Faculté des Sciences Jean Perrin, Université d’Artois, rue Jean Souvraz SP 18, 62307 Lens Cedex, FRANCE

currently: Institute for Advanced Study, Einstein Dr., Princeton, NJ 08540, USA

  E-mail address: karpenko@euler.univ-artois.fr, karpenko@ias.edu